### v1

# Sparse Voxels Rasterization: Real-time High-fidelity Radiance Field Rendering

Abstract
We propose an efficient radiance field rendering algorithm that incorporates a rasterization process on sparse voxels without neural networks or 3D Gaussians. There are two key contributions coupled with the proposed system. The first is to render sparse voxels in the correct depth order along pixel rays by using dynamic Morton ordering. This avoids the well-known popping artifact found in Gaussian splatting. Second, we adaptively fit sparse voxels to different levels of detail within scenes, faithfully reproducing scene details while achieving high rendering frame rates. Our method improves the previous neural-free voxel grid representation by over 4db PSNR and more than 10x rendering FPS speedup, achieving state-of-the-art comparable novel-view synthesis results. Additionally, our neural-free sparse voxels are seamlessly compatible with grid-based 3D processing algorithms. We achieve promising mesh reconstruction accuracy by integrating TSDF-Fusion and Marching Cubes into our sparse grid system.
Figure 1: We propose SVRaster, a novel framework for multi-view reconstruction and novel view synthesis. (a) Sparse voxel representation effectively captures the volume density and radiance field of the scene, without the need for neural networks or sparse-points prior initialization. (b) Using our customized sparse voxel rasterizer, we can learn the underlying 3D scene efficiently and achieve state-of-the-art performance in both rendering quality and speed. (c) Notably, our sparse voxel representation can be naturally integrated with classic 3D algorithms such as TSDF-Fusion [7, 31, 8] and Marching Cubes [23], making it flexible and suitable for a wide range of applications.
1 Introduction
Gaussian splatting [19] has emerged as one of the most promising solutions for novel view synthesis. It has drawn wide attention across multiple communities due to its exceptional rendering speed and ability to capture the nuanced details of a scene. Nevertheless, Gaussian splatting in its base form has two key limitations: first, sorting Gaussians based on their centers does not guarantee proper depth ordering. It may result in popping artifacts [32] (i.e., sudden color changes for consistent geometry) when changing views. Second, the volume density of a 3D point is ill-defined when covered by multiple Gaussians. This ambiguity makes surface reconstruction non-trivial.
On the other hand, grid representations inherently avoid these issues—both ordering and volume are well-defined. However, due to ray casting, the rendering speed of these methods [37, 5, 4, 38, 45, 22] is slower than that of Gaussian Splatting. This raises the question: is it possible to take the best of both worlds? Can we combine the efficiency of Gaussian splatting with the well-defined grid properties?
Fortunately, the answer is yes. Our key insight is to revisit voxels—a well-established primitive with decades of history. Voxel representations are inherently compatible with modern graphics engines and can be rasterized efficiently. Additionally, previous work has demonstrated their ability to model scene volume densities, despite through volume ray casting. This positions voxels as the perfect bridge between rasterization and volumetric representation. However, naively adopting voxels does not work well [11]. Since a scene may consist of different levels of detail.
With these observations in mind, we present SVRaster, a novel framework that combines the efficiency of rasterization in 3DGS with the structured volumetric approach of grid-based representations. SVRaster leverages (1) multi-level sparse voxels to model 3D scenes and (2) implements a dynamic Morton order encoding that facilitates the rasterization rendering from our adaptive-sized sparse voxel representation. Our rendering is free from popping artifacts because the 3D space is partitioned into disjoint voxels, and our sorting ensures the correct rendering order. Moreover, thanks to the volumetric nature and neural-free representation of SVRaster, our sparse voxel grid can be easily and seamlessly integrated with classical grid-based 3D processing algorithms, as demonstrated with the incorporated sparse-voxel TSDF Fusion and Marching Cubes.
We show that our method is training fast, rendering fast, and achieves novel-view synthesis quality comparable to the state-of-the-art. We extract meshes by the integrated sparse-voxel TSDF-Fusion and sparse-voxel Marching Cubes from the trained sparse voxel grids. Results on mesh reconstruction benchmarks show our good balance between surface accuracy and processing speed.
2 Related Work
Differentiable volume rendering has made significant strides in 3D scene reconstruction and novel-view synthesis tasks. Neural Radiance Fields (NeRF) [27] laid the foundation for this progress by optimizing a volumetric function, parameterized by multi-layer perceptrons (MLPs) to encode both geometry and appearance through differential rendering. Subsequent studies focus on accelerating speed by decomposing large MLPs into grid-based representations, with a shallow MLP typically still being employed. Several grid representations have been explored, including dense grids [37], sparse grids [11], factorized grids [5], tri-planes [4], and hash grids [29]. The most similar scene representation to ours is the sparse grid by Plenoxels [11]. We both explicitly model 3D scenes into voxels without neural networks. Our quality and scalability are significantly improved due to adaptive and non-uniform voxels, while our rendering is much faster because of rasterization. Another scalability issue of the previous grid-based methods is that they still use some sort of dense 3D grid. For instance, dense occupancy grids [37, 5, 29, 21] for free-space skipping or dense pointer grid [11] to support sparse voxel lookup. We do not use any 3D dense data structures.
3D Gaussian Splatting (3DGS) [19] takes a different approach by representing scenes with 3D Gaussian primitives and using rasterization for rendering, achieving state-of-the-art trade-off for quality and speed. Our work is inspired by the efficiency of using a rasterizer by 3DGS. However, 3DGS exhibits view-inconsistent popping artifacts due to inaccurate rendering order and primitive overlapping. Some recent works mitigate this artifact [32, 28], but completely resolving it significantly harms rendering speed [24]. Our method does not suffer from the popping artifact. Typically, 3DGS [19] initializes Gaussians from the triangulated points via COLMAP [35], which is especially critical for unbounded scenes. The sparse points geometry is also employed to guide the training of volumetric representations and has shown improvement in some NeRF-based methods [9, 12]. In this work, we do not use the sparse points.
Mesh reconstruction is an important extended topic for both volumetric-based and GS-based methods. Recent 3DGS variants [16, 14] with proper regularization have shown a good speed-accuracy trade-off. However, directly extracting meshes from 3DGS remains challenging, so these methods still rely on volumetric-based post-processing algorithms to extract meshes. NeRF ariants [41, 46, 38, 22, 45, 42] improve surface quality by rendering a signed distance field (SDF) instead of density field. The well-defined volume makes it straightforward to extract a mesh from the isosurface. Our sparse voxels are also a volumetric representation. Though our voxel also follows volumetric representation focusing on the density field, we still can achieve promising accuracy with minimal time.
For efficient rendering, another line of research is to convert (or distill) the geometry and appearance of a high-quality but slow model into other efficient kind of representations such as smaller MLPs [33, 40], meshes [43, 47, 34], grids [10], or the recent Gaussians [30]. Typically, the powerful Zip-NeRF [3] is employed as the teacher model which requires hours of training per scene. Note that using strong prior can be regarded as a different setup from ours.
Finally, the sparse voxel is also a popular representation for 3D processing and understanding [6, 13, 39, 44]. We mainly focus on fast and high-quality rendering while scene understanding is out of the scope of our work.
3 Approach
We present our approach as follows. First in Sec. 3.1, we introduce our sparse voxels scene representation and our rasterizer for rendering sparse voxels into pixels. Later in Sec. 3.2, we describe the progressive scene optimization strategy, which is designed for faithfully reconstructing the scene from multi-view images using our sparse voxels.
3.1 Sparse Voxels Rasterization
Recent neural radiance field rendering approaches, such as NeRF [27] and 3DGS [19] variants, use the following alpha composition equation to render a pixel’s color $\mathrm{C}$:
$\mathrm{C}$ $=\sum\nolimits_{i=1}^{N}T_{i}\cdot\alpha_{i}\cdot c_{i}~{},$ $T_{i}$ $=\prod\nolimits_{j=1}^{i-1}\left(1-\alpha_{j}\right)~{},$ (1)
where $\alpha_{i}\in\mathbb{R}_{\in[0,1]}$ and $c_{i}\in\mathbb{R}_{\in[0,1]}^{3}$ are alpha and view-dependent color at the $i$-th sampled point or primitive on the pixel-ray. The quantity $T_{i}$ is known as the transmittance. At a high-level, the difference between each method comes down to: i) how they find the $N$ points or primitives to composite a pixel-ray and ii) how the rgb and alpha values are determined from their scene representations.
Our sparse-voxel method follows the same principle. In Sec. 3.1.1, we provide details of our sparse voxels scene representation and how the $\alpha_{i}$ and $c_{i}$ in Eq. 1 are computed from a voxel. In Sec. 3.1.2, we present our rasterizer, that gathers the $N$ voxels to be composited for each pixel.
Figure 2: Sparse voxels scene representation. (Left) We allocate voxel under an Octree layout. Each voxels has its own Spherical Harmonic coefficient for view-dependent appearance. The color field is approximated as a constant inside a voxel when rendering a view for efficiency. The density field is trilienarly varied inside a voxel and is modeled by the density values on the corner grid points (i.e., the black dots $\bullet$) of each voxel. The grid points densities are shared between adjacent voxels. (Right) We evenly sample $K$ points inside the segment of ray-voxel intersection to compute volume integration for its alpha value contributing to the pixel ray. See Sec. 3.1.1 for details.
3.1.1 Scene Representation
We first describe the grid layout for allocating our sparse voxels and then derive the alpha value, view-dependent color, and other geometric properties needed for the composite rendering of a voxel.
Sparse voxel grid.
Our SVRaster constructs 3D scenes using a sparse voxel representation. We allocate voxels following an Octree space partition rule (i.e., Octree layout as illustrated in Fig. 2(a)) for two reasons necessary for achieving high-quality results. First, it facilitates the correct rendering order of voxels with various sizes (Sec. 3.1.2). Second, we can adaptively fit the sparse voxels to different scene level-of-details (Sec. 3.2). Note that our representation does not replicate a traditional Octree data structure with parent-child pointers or linear Octree. Specifically, we only keep voxels at the Octree leaf nodes without any ancestor nodes. Our sorting-based rasterizer will project voxels to image space and guarantee all voxels are in the correct order when rendering. In sum, we store individual voxels in arbitrary order without the need to maintain a more complex data structure, thanks to the flexibility provided by our rasterizer.
We choose a maximum level of detail $L$ ($=16$ in this work) that defines a maximum grid resolution at $(2^{L})^{3}$. Let $\mathbf{w}_{\mathrm{s}}\in\mathbb{R}$ be the Octree size and $\mathbf{w}_{\mathrm{c}}\in\mathbb{R}^{3}$ be the Octree center in the world space. The voxel index $v\!=\!\{i,j,k\}\in[0,\ldots,2^{L}\!\!\!-\!\!1]^{3}$ together with a Octree level $l\in[1,L]$ ($l=0$ represent root node and is not used) define voxel size $\mathbf{v}_{\mathrm{s}}$ and voxel center $\mathbf{v}_{\mathrm{c}}$ as:
$\mathbf{v}_{\mathrm{s}}$ $=\mathbf{w}_{\mathrm{s}}\cdot 2^{-l}~{},$ $\mathbf{v}_{\mathrm{c}}$ $=\mathbf{w}_{\mathrm{c}}-0.5\cdot\mathbf{w}_{\mathrm{s}}+\mathbf{v}_{\mathrm{s}}\cdot v~{}.$ (2)
Internally, we map the grid index to its Morton code using a well-known bit interleaving operation in the low-level CUDA implementation. Please see supplementary materials for more details.
Voxel alpha from density field.
Next, we present details for the geometry and appearance modeling of each voxel primitive. For scene geometry, we use eight parameters corresponding to the voxel corners to model a trilinear density field inside each voxel, denoted as $\mathbf{v}_{\mathrm{geo}}\in\mathbb{R}^{2\times 2\times 2}$. Sharing corners among adjacent voxels results in a continuous density field.
We also need an activation function to ensure a non-negative density value for the raw density from $\mathbf{v}_{\mathrm{geo}}$. For this purpose, we use exponential-linear activation:
$\mathrm{explin}(x)=\begin{cases}x&\text{if }x>1.1\\ \exp\left(\frac{x}{1.1}-1+\ln 1.1\right)&\text{otherwise}\end{cases}~{},$ (3)
which approximates $\mathrm{softplus}$ but is more efficient to compute. For a sharp density field inside a voxel, we apply the non-linear activation after trilinear interpolation [37, 18].
To derive the alpha value of a voxel contributing to the alpha composition formulation in Eq. 1, we evenly sample $K$ points in the ray segment of ray-voxel intersection as depicted in Fig. 2(b). The equation follows the numerical integration for volume rendering as in NeRF [27, 25]:
$\alpha=1-\exp\left(-\frac{l}{K}\sum_{k=1}^{K}\mathrm{explin}\left(\mathrm{interp}\left(\mathbf{v}_{\mathrm{geo}},\mathbf{q}_{k}\right)\right)\right)~{},$ (4)
where $l$ is the ray segment length, $\mathbf{q}_{k}$ is the local voxel coordinate of the $k$-th sample point, and $\mathrm{interp}(\cdot)$ indicates trilinear interpolation.
Voxel view-dependent color from SHs.
To model view-dependent scene appearance, we use $N_{\mathrm{shd}}$ degree SH. For increased efficiency, we assume the SH coefficients stay constant inside a voxel, denoted as $\mathbf{v}_{\mathrm{sh}}\in\mathbb{R}^{(N_{\mathrm{shd}}+1)^{2}\times 3}$. We approximate voxel colors as a function of the direction from the camera position $\mathbf{r}_{\mathrm{o}}$ to the voxel center $\mathbf{v}_{\mathrm{c}}$ instead of individual ray direction $\mathbf{r}_{\mathrm{d}}$ for the sake of efficiency following 3DGS:
$c=\max(0,\mathrm{sh\_eval}(\mathbf{v}_{\mathrm{sh}},\mathrm{normalize}(\mathbf{v}_{\mathrm{c}}-\mathbf{r}_{\mathrm{o}})))~{},$ (5)
which is the view-dependent color intensity of the voxel contributing to the pixel composition Eq. 1. Due to the approximation, the resulting SH color of a voxel can be shared by all covered pixels in the image rather than evaluating the SH for each intersecting ray.
Voxel normal.
The rendering of other features or properties is similar to rendering a color image by replacing the color term $c$ in Eq. 1 with the target modality like the normal of a voxel density field. For rendering efficiency, we assume the normal stays constant inside a voxel, which is represented by the analytical gradient of the density field at the voxel center:
$\vec{\bf n}=\mathrm{normalize}\left(\nabla_{\mathbf{q}}~{}\mathrm{interp}\left(\mathbf{v}_{\mathrm{geo}},\mathbf{q}_{\mathrm{c}}\right)\right)~{},$ (6)
where $\mathbf{q}_{\mathrm{c}}=(0.5,0.5,0.5)$ and the closed-form equations for forward and backward passes are in the supplementary material. Similar to the SH colors, the differentiable voxel normals are computed once in pre-processing and are shared by all the intersecting rays in the image.
Voxel depth.
Unlike colors and normals, the point depths to composite is efficient to compute so we do the same $K$ points sampling as in the voxel alpha value in Eq. 4 for a more precise depth rendering. We manually expand and simplify the forward and backward computation for small number of $K$ in our CUDA implementation. Please refer to supplementary materials for details.
Figure 3: Rasterization procedure. Refer to Sec. 3.1.2 for details.
Figure 4: Illustration of the rendering order. (a) In both cases, the smaller voxels should be rendered first but they will arranged behind the larger voxels if using voxel centers or the nearest corners as the sorting order. (b) We show the four types of Morton order under the 2D world. The voxel rendering order under an Octree node is depend on which world quadrant the ray direction is pointing to. (c) An toy example of sorting the Morton order encoding. All the ray directions going toward the up-right quadrant can use the sorted voxels for a correct rendering order. See Sec. 3.1.2 for details.
3.1.2 Rasterization Algorithm
The overview of our sparse voxel rasterization procedure is depicted in Fig. 3. We build our sorting-based rasterizer based on the highly efficient CUDA implementation of 3DGS [19]. The procedure is detailed in the following.
Projection to image space.
The first step of rasterization is projecting sparse voxels onto image space and assigning the voxels to the tiles (i.e., image patches) they cover. In practice, we project the eight corner points of each voxel. The voxel is assigned to all tiles overlapped with the axis-aligned bounding box formed by the projected eight points.
Pre-processing voxels.
For active voxels assigned to tiles of the target view, we gather their densities $\mathbf{v}_{\mathrm{geo}}$ from the grid points, compute view-dependent colors from their spherical harmonic coefficients with Eq. 5, and derive voxel normals by Eq. 6. The pre-processed voxel properties are shared among all pixels during rendering.
Sorting voxels.
For accurate rasterization, primitive rendering order is important. Similar to the challenge in 3DGS [19], using primitive centers or their closest distance to the camera can produce incorrect ordering, producing popping artifacts [32]. We show two incorrect ordering results using naive sorting criteria in Fig. 4(a). Thanks to the Octree layout (Sec. 3.1.1), we can sort by Morton order using our sparse voxels representation. As illustrated in Fig. 4(b), we can follow certain types of Morton order to render the voxels under an Octree node for correct ordering. The type of Morton order to follow is solely dependent on the positive/negative signs of the ray direction (the ray origin doesn’t matter). That is to say, we have eight permutations of Morton order for different ray directions in the 3D space. Finally, the generalization to multi-level voxels can be proved by induction. An ordering example in 2D with three levels is depicted in Fig. 4(c).
The sorting is applied for each image tile. In case all the pixels in a tile share the same ray direction signs, we can simply sort the assigned voxels by their type of Morton order. We handle the corner case when multiple Morton orders are required in supplementary materials.
Rendering pixels.
Finally, we proceed with alpha composition, Eq. 1, to render pixels. In our case, the $N$ primitives blend of a pixel-ray depends on the number of sparse voxels assigned to the tile that the pixel-ray belongs to. The computation of the alpha, color, and other geometric properties from our sparse voxels are described in Sec. 3.1.1. When rendering sparse voxels for a pixel-ray, we compute ray-aabb intersection to determine the ray segment to sample (for voxel alpha in Eq. 4) and skip some non-intersected sparse voxels. We do early termination of the alpha composition if the transmittance of a sparse voxel is below a threshold $T_{i}{<}h_{\mathrm{T}}$.
Anti-aliasing.
To mitigate aliasing artifacts, we render in $h_{\mathrm{ss}}$ times higher resolution and then apply image downsampling to the target resolution with an anti-aliasing filter.
3.2 Progressive Sparse Voxels Optimization
In this section, we describe the procedure to optimize a 3D scene from the input frames with known camera parameters using our SVRaster presented in Sec. 3.1.
Voxel max sampling rate.
We first define the maximum sampling rate $\mathbf{v}_{\mathrm{rate}}$ of each voxel on the training images, which reflects the image region a voxel can cover. A smaller $\mathbf{v}_{\mathrm{rate}}$ indicates that the voxel is more prone to overfitting due to less observation. We use $\mathbf{v}_{\mathrm{rate}}$ in our voxel initialization and subdivision process described later. Given $N_{\mathrm{cam}}$ training cameras, we estimate the maximum sampling rate of a voxel as follows with visualization in Fig. 5(a):
$\mathbf{v}_{\mathrm{rate}}$ $=\max\nolimits_{i}^{N_{\mathrm{cam}}}~{}\frac{\mathbf{v}_{\mathrm{s}}}{\mathbf{v}_{\mathrm{interval}}^{(i)}}~{},$ (7a)
$\mathbf{v}_{\mathrm{interval}}^{(i)}$ $=\underbrace{(\mathbf{v}_{\mathrm{c}}-\mathbf{r}_{\mathrm{o}}^{(i)})^{\intercal}\ell^{(i)}}_{\text{Voxel z-distance}}\cdot\underbrace{\frac{\tan\left(0.5\cdot\theta_{\text{fov-x}}^{(i)}\right)}{0.5\cdot W^{(i)}}}_{\text{Unit-distance pixel size}}~{},$ (7b)
where $\ell$ is camera lookat vector, $\theta_{\text{fov-x}}$ is camera horizontal field of view, and $W$ is image width. The sampling rate indicates the estimated number of rays along the image’s horizontal axis direction that may hit the voxel.
Figure 5: Visualization of voxel sampling rate and grid layout initialization. (a) We visualize the voxel sampling rate defined in Eq. 7. (b) We depict the foreground main region and the background region under different shell levels. In unbounded scenes, we apply different grid layout initialization strategies for foreground and background regions. See Sec. 3.2.1 for details.
3.2.1 Scene Initialization
Without employing an additional prior, we initialize all the parameters to constant. We start with volume density approaching zero by setting voxel raw density to a negative number $h_{\mathrm{geo}}$ so that the initial activated density $\mathrm{explin}(h_{\mathrm{geo}}){\approx}0$. We set the SH coefficients to zero for non-zero degrees and set the view-independent zero-degree component to yield gray color (i.e., intensity equal $0.5$). We detail the Octree grid layout initialization in the following.
Bounded scenes.
In case the scenes or the objects to reconstruct are enclosed in a known bounded region, we simply initialize the layout as a dense grid with $h_{\mathrm{lv}}$ levels and remove voxels unobserved by any training images. The number of voxels is $\leq(2^{h_{\mathrm{lv}}})^{3}$ after initialization.
Unbounded scenes.
For the unbounded scenes, we first split the space into the main and the unbounded background regions, depicted in Fig. 5(b), each with a different heuristic. We use the training camera positions to determine a cuboid for the main region. The cuboid center is set to the average camera positions and the radius is set to the median distance between the cuboid center and the cameras. The same as the bounded scenes, we initialize a dense grid with $h_{\mathrm{lv}}$ levels for the main region. For the background region, we allocate $h_{\mathrm{out}}$ level of background shells enclosing the main region, which means that the radius of the entire scene is $2^{h_{\mathrm{out}}}$ of the main region. In each background shell level, we start with the coarsest voxel size, i.e., $564^{3}{-}2^{3}{=}$ 56 voxels in each shell level. We then iteratively subdivide shell voxels with the highest sampling rate and remove voxels unobserved by any training cameras. The process repeats until the ratio of the number of voxels in the background and the main regions is $h_{\mathrm{ratio}}$. The number of voxels is $\leq(1+h_{\mathrm{ratio}})(2^{h_{\mathrm{lv}}})^{3}$ after initialization.
3.2.2 Adaptive Pruning and Subdivision
The initialized grid layout only coarsely covers the entire scene that should be adaptively aligned to different levels-of-detail for the scene during the training progress. We apply the following two operations every $h_{\mathrm{every}}$ training iterations to achieve this purpose.
Pruning.
We compute the maximum blending weight ($T_{i}\alpha_{i}$) from Eq. 1 of each voxel using all the training cameras. We remove voxels with maximum blending weight lower than $h_{\mathrm{prune}}$.
Subdivision.
Our heuristic is that a voxel with a larger training loss gradient indicates that the voxel region requires finer voxels to model. Specifically, we accumulate the subdivision priority as the following:
$\mathbf{v}_{\mathrm{priority}}=\sum\nolimits_{\mathbf{r}\in R}\left\|\alpha(\mathbf{r})\cdot\partialderivative{\mathcal{L}(\mathbf{r})}{\alpha(\mathbf{r})}\right\|~{},$ (8)
where $R$ is the set of all training pixel rays throughout the $h_{\mathrm{every}}$ iterations and $\mathcal{L}(\mathbf{r})$ is the training loss of the ray. The gradient is weighted by alpha values contributed from the voxel to the ray. Higher $\mathbf{v}_{\mathrm{priority}}$ indicates higher subdivision priority. To prevent voxels from overfitting few pixels, we set the priority to zero for voxels with maximum sampling rate lower than a sampling rate threshold $\mathbf{v}_{\mathrm{rate}}<2h_{\mathrm{rate}}$. Finally, we select the voxels with priority above the top $h_{\mathrm{percent}}$ percent to subdivide, i.e., the total number of voxels is increased by $(h_{\mathrm{percent}}\cdot(8-1))$ percent. Note that we only keep the leaf nodes in the Octree layout so we remove the source voxels once they are subdivided.
When voxels are pruned and subdivided, the voxel Spherical Harmonic (SH) coefficients and the grid point densities need to be updated accordingly. The SH coefficients are simply pruned together with voxels and duplicated to the subdivided children voxels. Grid point densities are slightly more complex as the eight voxel corner grid points are shared between adjacent voxels (Fig. 2). We remove a grid point only when it does not belong to any voxel corners. When subdividing, we use trilinear interpolation to compute the densities of the new grid points. The duplicated grid points are merged and their densities are averaged.
3.2.3 Optimization objectives
We use MSE and SSIM as the photometric loss between the rendered and the ground truth images. The overall training objective is summarized as:
$\mathcal{L}=\mathcal{L}_{\mathrm{mse}}+\lambda_{\mathrm{ssim}}\mathcal{L}_{\mathrm{ssim}}\\ +\lambda_{\mathrm{T}}\mathcal{L}_{\mathrm{T}}+\lambda_{\mathrm{dist}}\mathcal{L}_{\mathrm{dist}}+\lambda_{\mathrm{R}}\mathcal{L}_{\mathrm{R}}+\lambda_{\mathrm{tv}}\mathcal{L}_{\mathrm{tv}}~{},$ (9)
where $\lambda$ are the loss weights, $\mathcal{L}_{\mathrm{T}}$ encourages the final ray transmittances to be either zero or one, $\mathcal{L}_{\mathrm{dist}}$ is the distortion loss [2], $\mathcal{L}_{\mathrm{R}}$ is the per-point rgb loss [37], and $\mathcal{L}_{\mathrm{tv}}$ is the total variation loss on the sparse density grid. In mesh extraction task, we also add the depth-normal consistency loss from 2DGS [16]:
$\mathcal{L}_{\mathrm{mesh}}=\lambda_{\mathrm{n\text{-}dmean}}\mathcal{L}_{\mathrm{n\text{-}dmean}}+\lambda_{\mathrm{n\text{-}dmed}}\mathcal{L}_{\mathrm{n\text{-}dmed}}~{},$ (10)
where both losses encourage alignment between rendered normals and depth-derived normals from mean and median depth. More details are in the supplementary materials.
3.2.4 Sparse-voxel TSDF Fusion and Marching Cubes
Our sparse voxels can be seamlessly integrated with grid-based algorithms. To extract a mesh, we implement Marching Cubes [23] to extract the triangles of an isosurface over density from the sparse voxels. The duplicated vertices from adjacent voxels are merged to produce a unique set of vertices. When the adjacent voxels belong to different Octree levels, the extracted triangle may not be connected as the density field is not continuous for voxels in different levels. Such discontinuities can be removed by simply subdividing all voxels to the finest level.
Deciding the target level set for extracting the isosurface can be tricky for the density field. Instead, we implement sparse-voxel TSDF-Fusion [7, 31, 8] to compute the truncated signed distance values of the sparse grid points. We can then directly extract the surface of the zero-level set using the former sparse-voxel Marching Cubes. Future extensions of our method could directly model signed distance fields following NeuS [41] and VolSDF [46]. The sparse-voxel TSDF-Fusion can still be beneficial to directly initialize our sparse voxel representation from sensor depth.
Figure 6: A qualitative comparison with 3DGS [19]. Our result here corresponding to the base version in Tab. 1. We achieve similar visual quality comparing to 3DGS. Note that 3DGS use the coarse geometry from SfM while we do not rely on this prior.
Figure 7: Visualization of the reconstructed surface. We show the rendering images, normal maps, and the final meshes on Tanks&Temples and DTU datasets. Note that we only model the scene with density field and do not use the coarse geometry prior from SfM sparse points in this work.
[…]
Plenoxels [11] $<$10 $\sim$30m 0.463 23.08 0.626
INGP [29] $\sim$10 $\sim$5m 0.302 25.68 0.705
3DGS [19] 131 24m 0.216 27.45 0.815
Ours-small 240 8m 0.213 26.78 0.796
Ours-base 137 12m 0.197 27.10 0.807
Ours-large 92 17m 0.192 27.38 0.814
Re-evaluate on our machine using the public code.
Table 1: Novel-view synthesis results comparison on Mip-NeRF360 dataset [2]. The results are averaged from 4 indoor scenes and 5 outdoor scenes.
3DGS variants Ours
Method 3DGS [19] StopThePop [32] 3DGRT [28] EVER [24] base large
Rasterization
No popping $\triangle$ $\triangle$
[…]
† Re-evaluated on our machine using the public code.
‡ We scale the FPS to align their reported 3DGS FPS to our reproduced 3DGS.
Table 2: Comparison with recent 3DGS variants on Mip-NeRF360 dataset [2]. The LPIPS values here are evaluated with the correct intensity scale between $[-1,1]$ following EVER [24]. 3DGRT and EVER use ray tracing approach instead of rasterization, where only the concurrent EVER can totally resolve the popping artifact of GS-based rendering. Please find supplementary for more visual comparison.
Tanks&Temples Deep Blending
[…]
Plenoxels [11] 13 0.379 21.08 11 0.510 23.06
INGP [29] 14 0.305 21.92 3 0.390 24.96
M-NeRF360 [2] $<$1 0.257 22.22 $<$1 0.245 29.40
3DGS [19] 180 0.176 23.75 140 0.244 29.60
Ours 110 0.165 22.73 260 0.243 29.57
Re-evaluate on our machine using the public code.
Table 3: Comparison on Tanks&Temples [20] and Deep Blending [15] datasets. We follow 3DGS [19] to use two outdoor scenes from Tanks&Temples and two indoor scenes from Deep Blending.
Figure 8: Failuare case. On scenes with severe exposure variation of training views, our method struggles and produces clear boundary of different brightnesses and allocates many floaters. 3DGS on the other hand is less sensitive to photometric variation of GT. This explains our worse PSNR and FPS on Tanks&Temples in Tab. 3.
4 Experiments
4.1 Implementation Details
We use the following implementation details except stated otherwise. We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels (Secs. 3.2.1 and 5). The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.1$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for mesh reconstruction task. We train our model for 20K iterations. We find the merit of training longer is marginal to us. The voxels are pruned and subdivided every $h_{\mathrm{every}}{=}1{,}000$ iterations until 15K iterations. We prune voxels with less than $h_{\mathrm{prune}}{=}0.05$ maximum blending weights and subdivide top $h_{\mathrm{percent}}{=}5$ percent of voxels with higher priority (Eq. 8). More details are provided in the supplement.
We use the standard LPIPS [48], SSIM, and PSNR metrics to evaluate novel-view quality. We directly employ the LPIPS implementation from prior work [19], which uses $[0,1]$ image intensity range with VGG [36] network. For mesh accuracy, we follow the benchmarks to use F-score and chamfer distance. We use the test-set images to measure the FPS on a desktop computer with a 3090 Ti GPU.
4.2 Novel-view Synthesis
We show results on the real-world datasets in this section. Results on synthetic data are in supplementary materials.
In Tab. 1, we show the quantitative comparison on the MipNeRF-360 dataset. We mainly compare the representative scene representations with efficient training and rendering. Regarding frame rates, 3DGS [19] and ours with rasterization achieve much higher frame rates compared to the previous ray casting and sampling approach [11, 29]. Zip-NeRF [3] still achieves the best novel-view quality but is prohibitively slow to train and render. Compared to the competitive 3DGS [19], our base setup achieves similar FPS, better LPIPS, and training time, but worse in PSNR and SSIM. This may be explained by the qualitative comparison in Fig. 6. We find 3DGS rendering is typically smoother than ours due to its Gaussian representation while 3DGS sometimes fails to recover finer details. Some other important factors include: i) 3DGS uses the SfM sparse points to initialize while we do not and ii) we use different strategies to progressively grow our model. However, it is difficult to align everything for a side-by-side comparison due to the difference in our scene representations.
We also train a small variant by pruning more voxels with $h_{\mathrm{prune}}{=}0.15$ and a large variant by pruning less $h_{\mathrm{prune}}{=}0.01$, higher supersample rate $h_{\mathrm{ss}}{=}1.5$, and $K{=}2$ sample points inside each voxel. Our small variant achieves similar LPIPS with +80% FPS compared to 3DGS. Our large variant has 92 FPS with LPIPS close to Zip-NeRF.
More comparisons with recent 3DGS variants are in Tab. 2. 3DGS [19] has popping artifacts due to the ordering and overlapping issues. StopThePop [32] uses running sort and 3DGRT [28] uses ray tracing for accurate ordering but drops FPS by 28% and 67% respectively. EVER [24] further handles the Gaussian overlapping cases but with even less FPS. Our method ensures correct ordering (Sec. 3.1.2) with FPS and quality on par with the original 3DGS.
We also compare results on two more datasets in Tab. 3. On the Tanks&Temples dataset [20], despite having better LPIPS, our PSNR, and FPS are much worse than 3DGS. The failure results are shown in Fig. 8. We notice that the training views of the Tanks&Temples dataset have large exposure variation. As a result, our method produces a clear brightness boundary in the scene with many floaters that slow down our rendering. On the Deep Blending [15] dataset, we achieve similar LPIPS and PSNR as 3DGS but with 85% higher FPS. The speed and quality of 3DGS and our method are scene-dependent. More results and discussion are in the supplementary.
Resolution of main $256^{3}$ $512^{3}$ $1024^{3}$ adaptive Plenoxels $640^{3}$
[…]
Table 4: Ablation experiments of adaptive and uniform voxel sizes. The results are evaluated on the indoor bonsai and the outdoor bicycle scenes from MipNeRF-360 dataset. The resolutions at the first row indicate the final grid resolution of the main foreground cuboid. Plenxoels is the previous fully-explicit voxel grid approach. Please refer to Sec. 4.3 for more discussion.
4.3 Ablation Studies
The rasterization process improves the rendering speed of sparse voxels, while the key to achieve high-quality results is the adaptive voxel size for different levels of detail. We show an ablation experiment in Tab. 4. Instead of subdividing voxels adaptively, the uniform voxel size variant subdivides all the voxels at certain training iterations until the grid resolution of the foreground main region reaches $256^{3}$, $512^{3}$, or $1024^{3}$. We also align the background and the pruning setup where the details are deferred to the supplementary. As shown in the table, the quality of uniform voxel size in $512^{3}$ resolution is much worse than the adaptive voxel size under similar FPS. Our machine with 24GB GPU memory fails to reach $1024^{3}$ grid resolution despite the voxels being pruned and sparse. Plenoxels [11] renders sparse voxels by ray casting and sampling following NeRF. Their FPS is significantly lower than our rasterization-based sparse voxel rendering. As they use a dense 3D pointer grid to support point sampling on rays, the model scalability for the unbounded background region is thus limited. To workaround, they apply a multi-sphere image with 64 layers to model the background. Conversely, we set the scene size as 32x larger ($h_{\mathrm{out}}{=}5$) than the foreground cuboid and direct model the entire scene by our sparse voxels. As a result, our uniform size variant with $512^{3}$ foreground resolution still outperforms Plenoxels with $640^{3}$ grid.
In our rasterizer, we use dynamic Morton order to ensure the sorting of voxels from different levels is correct. However, like the case in 3DGS, the popping artifact by the incorrect order is not a major factor in the numerical results. Instead, we provide more visualization in the supplementary videos to show its effect. See more ablation studies in the supplementary materials.
Tansk&Temples DTU
Method Geom. Init. F-$\uparrow$ Tr. $\downarrow$ Cf.$\downarrow$ Tr. $\downarrow$
NeRF [27] Density Random - - 1.49 hrs
NeuS [41] SDF Sphere 0.38 hrs 0.84 hrs
Voxurf [45] SDF Sphere - - 0.76 15m
Neuralangelo [22] SDF Sphere 0.50 hrs 0.61 hrs
3DGS [19] Gaussians SfM pts. 0.19 14m 1.96 11m
SuGaR [14] Gaussians SfM pts. 0.09 $\sim$1h 1.33 $\sim$1h
2DGS [16] Gaussians SfM pts. 0.32 16m 0.80 11m
Ours Density Constant 0.40 11m 0.76 5m
Table 5: Mesh results comparison on the Tanks&Temples [20] and DTU [17] datasets. Our method with volume density and constant field initialization achieves good accuracy-time trade-off on the unbounded-level and object-level datasets.
4.4 Mesh Reconstruction
We use sparse-voxel TSDF-Fusion and Marching Cubes (Sec. 3.2.4) to extract mesh from the optimized volume density and sparse grid layout. The results comparison on the large-scale Tanks&Temples and the object-scale DTU datasets is provided in Tab. 5. Neuralangelo [22] is still state-of-the-art in terms of mesh accuracy but needs hours of optimization per scene. Our method achieves a good balance of accuracy and training time on both datasets. Interestingly, our method with density modeling still achieves on-par geometric accuracy with some other methods using a signed distance field (SDF). This shows that a density field with proper regularization and strong scene modeling capacity could still potentially achieve good accuracy on both large-scale and object-scale scenes. Compared to the recent GS-based methods, we do not use the coarse geometry estimated by SfM and still achieve better accuracy and training time. Some qualitative examples are in Fig. 7. More results are in the supplementary materials.
5 Conclusion
This work presents a novel differentiable radiance field rendering system that integrates an efficient rasterizer with a multi-level sparse voxel scene representation. We reconstruct a scene from the input multi-view images by adaptively fitting the sparse voxels into different levels-of-detail. We also incorporate some classical grid-based 3D processing algorithms into our system including sparse-voxel TSDF-Fusion and Marching Cubes for mesh extraction. Experiments on novel-view synthesis and mesh reconstruction tasks demonstrate our effectiveness, achieving high FPS rendering, fast training, and high result quality.
Supplementary Material
We provide more details of our method and implementation in Secs. A, B and C. Some content overlaps with the main paper in the interest of being self-contained. More results and discussion are found in Secs. D and E.
A More Details of Our Representation
A.1 Details of Sparse Voxels Grid
Recall that our SVRaster allocates voxels following an Octree layout but does not replicate a traditional Octree data structure with parent-child pointers or linear Octree. We only keep voxels at the Octree leaf nodes without any ancestor nodes and store individual voxels in arbitrary order without the need to maintain a more complex data structure.
The maximum level of detail is set to $L{=}16$ that defines the finest grid resolution at $65536^{3}$. Note that this is only for our CUDA-level implementation convenience. We leave it as future work to extend to an arbitrary number of levels as we find that $16$ levels are adequate for the scenes we experimented with in this work.
Let $\mathbf{w}_{\mathrm{s}}\in\mathbb{R}$ be the Octree size and $\mathbf{w}_{\mathrm{c}}\in\mathbb{R}^{3}$ be the Octree center in the world space. The voxel index $v\!=\!\{i,j,k\}\in[0,\ldots,2^{L}\!\!\!-\!\!1]^{3}$ together with an Octree level $l\in[1,L]$ ($l=0$ represent root node and is not used) define voxel size $\mathbf{v}_{\mathrm{s}}$ and voxel center $\mathbf{v}_{\mathrm{c}}$ as:
$\mathbf{v}_{\mathrm{s}}$ $=\mathbf{w}_{\mathrm{s}}\cdot 2^{-l}~{},$ $\mathbf{v}_{\mathrm{c}}$ $=\mathbf{w}_{\mathrm{c}}-0.5\cdot\mathbf{w}_{\mathrm{s}}+\mathbf{v}_{\mathrm{s}}\cdot v~{}.$ (11)
Internally, we map the grid index to its Morton code by a well-known bit interleaving operation, which is helpful to implement our rasterizer detailed later. A Python pseudocode is provided in Sec. A.1. {listing}[t] {minted} [frame=single,fontsize=]python MAX_NUM_LEVELS = 16
def to_octpath(i, j, k, lv): # Input # (i,j,k): voxel index. # lv: Octree level. # Output # octpath: Morton code octpath: int = 0 for n in range(lv): bits = 4*(i&1) + 2*(j1) + (k1) octpath —= bits ¡¡ (3*n) i = i ¿¿ 1 j = j ¿¿ 1 k = k ¿¿ 1 octpath = octpath ¡¡ (3*(MAX_NUM_LEVELS-lv)) return octpath
def to_voxel_index(octpath, lv): # Input # octpath: Morton code # lv: Octree level. # Output # (i,j,k): voxel index. i: int = 0 j: int = 0 k: int = 0 octpath = octpath ¿¿ (3*(MAX_NUM_LEVELS-lv)) for n in range(lv): i —= ((octpath0b100)¿¿2) ¡¡ n j —= ((octpath0b010)¿¿1) ¡¡ n k —= ((octpath0b001)) ¡¡ n octpath = octpath ¿¿ 3 return (i, j, k) Pseudocode for conversion between voxel index and Morton code. See Sec. A.1 for details. {listing}[t] {minted} [frame=single,fontsize=]python def ray_aabb(vox_c, vox_s, ro, rd): # Input # vox_c: Voxel center position. # vox_s: Voxel size. # ro: Ray origin. # rd: Ray direction. # Output # a: Ray enter at (ro + a * rd). # b: Ray exit at (ro + b * rd). # valid: If ray hit the voxel. c0 = (vox_c - 0.5 * vox_s - ro) / rd c1 = (vox_c + 0.5 * vox_s - ro) / rd a = torch.minimum(c0, c1).max() b = torch.maximum(c0, c1).min() valid = (a ¡= b) & (a ¿ 0) return a, b, valid Pseudocode for intersecting ray and a axis-aligned voxel. See Sec. A.2 for details.
A.2 Details of Voxel Alpha from Density
A voxel density field is parameterized by eight parameters attached to its corners $\mathbf{v}_{\mathrm{geo}}\in\mathbb{R}^{2\times 2\times 2}$, which is denoted as $\mathbf{V}$ for brevity in the later equations. We use the exponential-linear activation function to map the raw density to non-negative volume density. We visualize exponential-linear and Softplus in Fig. 9. Exponential-linear is similar to Softplus but more efficient to compute on a GPU. For a sharp density field inside a voxel, we apply the non-linear activation after trilinear interpolation [37, 18].
We evenly sample $K$ points in the ray segment of ray-voxel intersection to derive the voxel alpha value contributing to the pixel ray. First, we compute the ray voxel intersection point by Sec. A.1, which yields the ray distances $a$ and $b$ for the entrance and exit points along the ray with ray origin $\mathbf{r}_{\mathrm{o}}\in\mathbb{R}^{3}$ and ray direction $\mathbf{r}_{\mathrm{d}}\in\mathbb{R}^{3}$. The coordinate of $k$-th of the K sample points is:
$t_{k}$ $=a+\frac{k-0.5}{K}\cdot(b-a)$ (12a)
$\mathbf{p}_{k}$ $=\mathbf{r}_{\mathrm{o}}+t_{k}\cdot\mathbf{r}_{\mathrm{d}}$ (12b)
$\mathbf{q}_{k}$ $=\left(\mathbf{p}_{k}-(\mathbf{v}_{\mathrm{c}}-0.5\cdot\mathbf{v}_{\mathrm{s}})\right)\cdot\frac{1}{\mathbf{v}_{\mathrm{s}}}~{},$ (12c)
where $\mathbf{p}_{k}{\in}\mathbb{R}^{3}$ is in the world coordinate and $\mathbf{q}_{k}{\in}\mathbb{R}_{\in[0,1]}^{3}$ is in the local voxel coordinate. The local coordinate $\mathbf{q}$ is used to sample voxel by trilinear interpolation:
$\mathrm{interp}(\mathbf{V},\mathbf{q})=\begin{bmatrix}{\scriptstyle(1-\mathbf{q}_{x})\cdot(1-\mathbf{q}_{y})\cdot(1-\mathbf{q}_{z})}\\ {\scriptstyle(1-\mathbf{q}_{x})\cdot(1-\mathbf{q}_{y})\cdot(\mathbf{q}_{z})}\\ {\scriptstyle(1-\mathbf{q}_{x})\cdot(\mathbf{q}_{y})\cdot(1-\mathbf{q}_{z})}\\ {\scriptstyle(1-\mathbf{q}_{x})\cdot(\mathbf{q}_{y})\cdot(\mathbf{q}_{z})}\\ {\scriptstyle(\mathbf{q}_{x})\cdot(1-\mathbf{q}_{y})\cdot(1-\mathbf{q}_{z})}\\ {\scriptstyle(\mathbf{q}_{x})\cdot(1-\mathbf{q}_{y})\cdot(\mathbf{q}_{z})}\\ {\scriptstyle(\mathbf{q}_{x})\cdot(\mathbf{q}_{y})\cdot(1-\mathbf{q}_{z})}\\ {\scriptstyle(\mathbf{q}_{x})\cdot(\mathbf{q}_{y})\cdot(\mathbf{q}_{z})}\end{bmatrix}^{\intercal}\begin{bmatrix}\mathbf{V}_{000}\\ \mathbf{V}_{001}\\ \mathbf{V}_{010}\\ \mathbf{V}_{011}\\ \mathbf{V}_{100}\\ \mathbf{V}_{101}\\ \mathbf{V}_{110}\\ \mathbf{V}_{111}\end{bmatrix}~{},$ (13)
where the subscript in this equation indicates the $x,y,z$ components of the vector $\mathbf{q}$ and the sample index is omitted. Following NeRF [27, 25], we use quadrature to compute the integrated volume density for alpha value:
$\alpha$ $=1-\exp\left(-\frac{l}{K}\sum_{k=1}^{K}\mathrm{explin}\left(v_{k}\right)\right)$ (14a)
$v_{k}$ $=\mathrm{interp}\left(\mathbf{V},\mathbf{q}_{k}\right)$ (14b)
$l$ $=(b-a)\cdot\|\mathbf{r}_{\mathrm{d}}\|~{},$ (14c)
where $l$ is the ray segment length. The gradient with respect to the voxel density parameters is:
[…]
Figure 9: Activation functions. We use exponential-linear activation to softly map raw density to non-negative volume density. Exp-lin activation is about two times faster to compute in CUDA, which is 21.5M operations per second in a CUDA thread comparing to 11.8M of Softplus. See Sec. A.2 for more details.
A.3 Details of Voxel Normal
Recall that we approximate the normal field as constant inside a voxel for efficiency, which is represented by the analytical gradient of the density field at the voxel center $\mathbf{q}^{\mathrm{(c)}}$. Thanks to the neural-free representation, we derive closed-form equations for forward and backward passes instead of relying on double backpropagation of autodiff. The unnormalized voxel normal in the forward pass is:
$\nabla_{\mathbf{q}}\mathrm{interp}(\mathbf{V},\mathbf{q}^{\mathrm{(c)}})=0.25~{}\cdot$
$\begin{bmatrix}{\scriptstyle(\mathbf{V}_{100}+\mathbf{V}_{101}+\mathbf{V}_{110}+\mathbf{V}_{111})-(\mathbf{V}_{000}+\mathbf{V}_{001}+\mathbf{V}_{010}+\mathbf{V}_{011})}\\ {\scriptstyle(\mathbf{V}_{010}+\mathbf{V}_{011}+\mathbf{V}_{110}+\mathbf{V}_{111})-(\mathbf{V}_{000}+\mathbf{V}_{001}+\mathbf{V}_{100}+\mathbf{V}_{101})}\\ {\scriptstyle(\mathbf{V}_{001}+\mathbf{V}_{011}+\mathbf{V}_{101}+\mathbf{V}_{111})-(\mathbf{V}_{000}+\mathbf{V}_{010}+\mathbf{V}_{100}+\mathbf{V}_{110})}\end{bmatrix}$ (16)
For the backward pass, the gradient with respect to a density parameter is:
$\nabla_{\mathbf{V}_{ijk}}\nabla_{\mathbf{q}}\mathrm{interp}(\mathbf{V},\mathbf{q}^{\mathrm{(c)}})=0.25\cdot\begin{bmatrix}2i-1\\ 2j-1\\ 2k-1\end{bmatrix}~{}.$ (17)
A.4 Details of Voxel Depth
Voxel depths are efficient to compute compared to view-dependent colors and normals so we do the same $K$ points sampling as in the voxel alpha value. Unlike colors and normals, which are approximated by constant inside each voxel, the depth values of each sample point inside a voxel are different so we need to incorporate the point depth in Eq. 12a into the local alpha composition in Eq. 14. Let
$\alpha_{k}=1-\exp\left(-\frac{l}{K}\cdot\mathrm{explin}\left({\scriptstyle\mathrm{interp}\left(\mathbf{V},\mathbf{q}_{k}\right)}\right)\right)$ (18)
be the alpha value of the $k$-th sampled point. The voxel local depth is:
$d=\sum_{k=1}^{K}\left(\prod_{j=1}^{k-1}(1-\alpha_{j})\right)\cdot\alpha_{k}\cdot t_{k}~{}.$ (19)
Finally, the pixel depth is composited by ${\scriptstyle\mathrm{D}{=}\sum_{i=1}^{N}T_{i}d_{i}}$ from the $N$ voxels, where $T_{i}$ is the ray transmittance when reaching the $i$-th voxel described in the main paper.
[…]
The voxel depth when $K{=}3$ is:
$d=\alpha_{1}t_{1}+(1-\alpha_{1})\alpha_{2}t_{2}+(1-\alpha_{1})(1-\alpha_{2})\alpha_{3}t_{3}~{}.$ (21)
The backward equations are:
[…]
B More Details of Voxel Rendering Order
Our sorting-based rasterizer is based on the efficient CUDA implementation done by 3DGS [19]. In the following, we first describe how the overall sorting pipeline works in Sec. B.1. We then dive more into the implementation of the dynamic Morton order in Sec. B.2 and its correctness proof in Sec. B.3.
A supplementary video is provided to show the effect of correct ordering and a few popping artifacts in comparison with 3DGS [19].
B.1 Overview
The goal in the sorting stage of the rasterizer is to arrange a list of voxels in near-to-far order for each image tile. To this end, 3DGS’s rasterizer duplicates a Gaussian for each image tile the Gaussian covers. A key-value pair is attached to each Gaussian duplication, where the tile index is assigned as the most significant bits of the sorting key. The bit field of the key-value pair is as follows:
key $=|\underbrace{\text{tile id}}_{\text{32 bits}}|\underbrace{\text{Gaussian z-depth}}_{\text{32 bits}}|$ (23a)
value $=|\underbrace{\text{Gaussian id}}_{\text{32 bits}}|$ (23b)
By doing so, all the duplicated Guassians assigned to the same image tile will be in the consecutive array segment after sorting with near-to-far z-depth ordering. In the later rendering stage, each pixel only iterates through the list of Gaussians of its tile for alpha composition.
In our case, we replace the primitive z-depth with a dynamic Morton order of voxels to ensure the rendering order is always correct. As there are eight different Morton orders to follow depending on the positive/negative signs of ray directions, dubbed ray sign bits, we further duplicate each voxel by the numbers of different ray sign bits it covers. The ray sign bits are also attached to each duplicated voxel. In the rendering stage, a pixel only composites voxels with the same attached ray sign bits when there are multiple ray sign bits in an image tile. Our bit field of the key-value pair is:
key $=|\underbrace{\text{tile id}}_{\text{16 bits}}|\underbrace{\text{dynamic Morton order}}_{\text{48 (=$3L$) bits}}|$ (24a)
value $=|\underbrace{\text{ray sign bits}}_{\text{3 bits}}|\underbrace{\text{voxel id}}_{\text{29 bits}}|$ (24b)
where $L{=}16$ is the maximum number of Octree levels. Note that the “voxel id” here is indexed to the 1D array location where we store the voxel. Not to be confused the grid $(i,j,k)$ index in Sec. A.1. The bit field arrangement is mainly for our implementation convenient to squeeze everything into 64 and 32 bits unsigned integers. In our current implementation, the maximum number of tiles is $655362^{16}{=}$ 65536, which is $4096{\times}4096$ maximum image resolution with $16{\times}16$ tile size; the maximum grid resolution is $(2^{16})^{3}{=}65536^{3}$; the maximum number of voxels is $2^{29}{\approx}500M$. We find this is more than enough for the scenes in our experiments. Future work can define custom data types with extra bits for GPU Radix sort [26] to increase the resolution limit. {listing}[t] {minted} [frame=single,fontsize=]python MAX_NUM_LEVELS = 16 order_tables = [ [0, 1, 2, 3, 4, 5, 6, 7], [1, 0, 3, 2, 5, 4, 7, 6], [2, 3, 0, 1, 6, 7, 4, 5], [3, 2, 1, 0, 7, 6, 5, 4], [4, 5, 6, 7, 0, 1, 2, 3], [5, 4, 7, 6, 1, 0, 3, 2], [6, 7, 4, 5, 2, 3, 0, 1], [7, 6, 5, 4, 3, 2, 1, 0], ]
def to_rd_signbits(rd): # Input # rd: Ray direction. # Output # signbits: Ray sign bits. return 4*(rd[0]¡0) + 2*(rd[1]¡0) + (rd[2]¡0)
def to_dynamic_morton_order(octpath, signbits): # Input # octpath: Voxel Octree Morton code. # signbits: The signbits the voxel care. # Output # order: The order for sorting. table = order_tables[signbits] order = 0 for i in range(MAX_NUM_LEVELS): order —= table[octpath & 0b111] ¡¡ (3*i) octpath = octpath ¿¿ 3 return order Pseudocode for dynamic Morton order. The mapping between voxel grid $(i,j,k)$ index and Octree Morton code octpath is detailed in Sec. A.1. In practice, the mapping from Octree Morton code to dynamic Morton order is done by a single bitwise $\mathrm{xor}$ operation instead of for-loop. More details in Sec. B.2.
B.2 Dynamic Morton Order
As described and illustrated in the main paper, there are eight types of Morton order to follow, each of which is for a certain type of positive/negative signs pattern of ray directions. We hard-code the eight types of Morton orders, which is used to remap every non-overlapping three bits (corresponding to different Octree levels) in the Octree Morton code of voxels (Sec. A.1):
${}b_{x}b_{y}b_{z}~{}a_{x}a_{y}a_{z}\mapsto...~{}f^{\mathrm{(k)}}(b_{x}b_{y}b_{z})~{}f^{\mathrm{(k)}}(a_{x}a_{y}a_{z})~{},$ (25)
where $f^{\mathrm{(k)}}:[0\cdots 7]\mapsto[0\cdots 7]$ is one of the eight permutation mappings. The pseudocode for computing the ray sign bits and the mapping function from Octree Morton code to dynamic Morton order is provided in Sec. B.1.
B.3 Proof of Correct Ordering
We prove the ordering correctness by induction. We focus on the case for (+,+,+)(+,+,+)( + , + , + ) ray directions. The proof can be generalized to the other types of ray direction signs by flipping the scene. The Morton order of the eight voxels in the first Octree level is illustrated in Fig. 10.
Recap that our sparse voxels only consist of the Octree leaf nodes without any ancestor nodes. Let $V^{\ell}$ be the space of all valid sparse voxel sets with maximum Octree level equal to $\ell$. Let $S(\ell)$ be the statement that:
“For all sparse voxel sets in $V^{\ell}$, their dynamic Morton order is always aligned with the near-to-far rendering order for all rays with (+,+,+)(+,+,+)( + , + , + ) direction signs.”
Base case.
When $\ell{=}1$, there is only one Octree level. The dynamic Morton order of the eight voxels for (+,+,+)(+,+,+)( + , + , + ) ray directions is illustrated in Fig. 10. The bit field from most to least significant bit is for $x$, $y$, and $z$ directions, respectively. As the ray is going toward $+x$ direction, we can always render the voxels in the $-x$ side ($000,001,010,011$) first before the voxels in the $+x$ side ($100,101,110,111$), which is aligned with the most significant bit of the Morton order. Similarly, for the voxels in the $-x$ side, we can render the voxels in the $-y$ side ($000,001$) before the $+y$ side ($010,011$) as the ray is going toward $+y$. Finally, we can see that the rendering order is correct if we iterate the voxels following the assigned Morton order for ray with (+,+,+)(+,+,+)( + , + , + ) directions.
Induction hypothesis.
Assume that $S(\ell)$ is true for some positive integer $\ell$.
Induction step.
We want to show $S(\ell)\implies S(\ell+1)$ is true. For any sparse voxel set $w\in V^{\ell+1}$, there exists a sparse voxel set $v\in V^{\ell}$ that can evolve into $w$ by: i) selecting a subset of voxels in $v$ to subdivide with the source voxels removed and ii) removing some of the voxels. The $S(\ell)$ indicates that the dynamic Morton order of $v$ has the correct rendering order. To extend for a new Octree level, three zero bits are first append to the least significant bit of the Morton order of every voxel in $v$, which does not affect the ordering. When subdividing a voxel, the eight child voxels share the same most significant $3\ell$ bits as the source voxel, while the least significant $3$ bits follow the same dynamic Morton order as in the base case Fig. 10. This reflects the fact that the new child voxels should keep the same relative order to the other voxels as their source parent voxels as the child voxels are all in the 3D space of the source voxels. The rendering ordering of the eight child voxels can also follow the same Morton order as the base case. That is the Morton order is still rendering-order correct after subdividing some voxels in $v$. Finally, removing voxels does not affect the ordering of the remaining others. In sum, the Morton order of $w$ also has the correct rendering order so $S(\ell)$ implies $S(\ell+1)$. By induction, $S(\ell)$ is true for all positive integer $\ell$.
Figure 10: Base case. Dynamic Morton order for (+,+,+)(+,+,+)( + , + , + ) ray direction signs under the base case with 1 Octree level. The three bits from left to right is for the $x$, $y$, and $z$ directions respectively. The rendering order is correct for all rays going toward (+,+,+)(+,+,+)( + , + , + ) direction. See Sec. B.3 for more details.
C Additional Implementation Details
We start the optimization from empty space with raw density set to $h_{\mathrm{geo}}{=}{-}10$. We use spherical harmonic (SH) with $N_{\mathrm{shd}}{=}3$ degrees. The learning rate is set to $0.05$ for the grid point densities, $0.01$ for zero-degree SH coefficients, and $0.00025$ for higher-degree SH coefficients. The momentum and the epsilon value of the Adam optimizer are set to $(0.1,0.99)$ and $1\text{e}{-}15$. The initial Octree level is $h_{\mathrm{lv}}{=}6$ (i.e., $64^{3}$ voxels) for the bounded scenes and the foreground main region of the unbounded scenes. To model unbounded scenes, we use $h_{\mathrm{out}}{=}5$ background shell levels with $h_{\mathrm{ratio}}{=}2$ times the number of foreground voxels. We use average frame color as the color coming from infinite far away for unbounded scenes. The early ray stopping threshold is set to $h_{\mathrm{T}}{=}1\mathrm{e}{-4}$ and the supersampling scale is set to $h_{\mathrm{ss}}{=}1.1$. Inside each voxel, we sample $K{=}1$ point for novel-view synthesis and $K{=}3$ points for the mesh reconstruction task.
We train our model for $20{,}000$ iterations. We find the benefit of training longer is marginal. The voxels are subdivided every $h_{\mathrm{every}}{=}1{,}000$ iterations until $15{,}000$ iterations, where the voxels with top $h_{\mathrm{percent}}{=}5$ percent priority are subdivided each time. We set $h_{\mathrm{rate}}{=}1$ and skip subdividing voxels with a maximum sampling rate below $2h_{\mathrm{rate}}$. The voxels are pruned every $h_{\mathrm{every}}{=}1{,}000$ iterations until $18{,}000$ iterations, where voxels with maximum blending weights less than a pruning threshold are removed. The pruning threshold is linearly increased from $0.0001$ at the first pruning to $h_{\mathrm{prune}}{=}0.05$ at the last pruning.
The loss weights are set to $\lambda_{\mathrm{ssim}}{=}0.02$, $\lambda_{\mathrm{T}}{=}0.01$, $\lambda_{\mathrm{dist}}{=}0.1$ after $10{,}000$ training iterations, $\lambda_{\mathrm{R}}{=}0.01$, $\lambda_{\mathrm{tv}}{=}1\text{e}{-}10$ until $10{,}000$ training iterations. For the mesh reconstruction task, the weights for normal-depth alignment self-consistency loss are set to $\lambda_{\mathrm{n\text{-}dmean}}{=}0.001$ and $\lambda_{\mathrm{n\text{-}dmed}}{=}0.001$ for mean and median depth respectively. The initial depths and normals are bad so the two normal-depth consistency loss is activated at the later training iterations. We find the median depth converges the fastest so we activate median depth-normal consistency loss at $3{,}000$ iterations, which also only regularizes the rendered depth as median depth is not differentiable. The mean depth-normal consistency loss is activated at $10{,}000$ iterations.
D Additional Ablation Studies
Resolution of main $256^{3}$ $512^{3}$ $1024^{3}$ adaptive
[…]
Table 6: Ablation experiments of adaptive and uniform voxel sizes. The resolutions at the first row indicate the final grid resolution of the main foreground cuboid. Note that OOM is abbreviation of the term, ‘out-of-memory’.
D.1 Novel-View Synthesis
We conduct comprehensive ablation experiments of our method using the indoor bonsai and the outdoor bicycle scenes from the MipNeRF-360 [2] dataset.
Adaptive voxel sizes.
In the main paper, we show that adaptive voxel size for different levels of detail is crucial to achieve high-quality results. The results are recapped in Tab. 6. We provide experiment details here. The starting point of the main foreground region is the same for all variants with $64^{3}$ dense voxels. Regarding the background region, using the same voxel size as the foreground region is impracticable for uniform-sized variants. Instead, each of the $5$ background shell voxels is uniformly subdivided by $4$ times as initialization for all the variants. The difference is that the uniform-sized variants subdivide all voxels each time until the grid resolution of the main region reaches $256^{3}$, $512^{3}$, or $1024^{3}$ instead of subdividing voxels adaptively as described in the main paper. The pruning setup remains the same for all variants. The result in Tab. 6 shows that adaptive voxel sizes are the key to solve the scalability issue of uniform-sized voxel, which achieves much better rendering quality with high render FPS.
More ablation studies for the hyperparameters.
We conduct more ablation experiments to show the effectiveness of the hyperparameters in Tabs. 9, 8, 8, 7, 10, 11, 12, 17, 15, 14, 13 and 16. We mark the adopted hyperparameter setup by “*” in the table rows. The setup of the marked rows across different tables can be different as we update the base setups in a rolling manner during the hyperparameter tuning stage. In each table, the other hyperparameter setups except the ablated one are the same. We discuss the experiments directly in the table captions to avoid the need for cross-referencing between the tables and the main text.
D.2 Mesh Reconstruction
To show the effectiveness of the mesh regularization losses, we use the Ignatius and the Truck scenes from TnT [20] dataset and three scans with id 24, 69, 122 from DTU [17] dataset for ablation studies. The results are shown in Tab. 18. While the normal-depth self-consistency losses do not improve novel-view synthesis quality in Tab. 17, the mesh accuracy is improved by an obvious margin with the regularizations.
[…]
no ss 111 13.5m 0.201 27.69 0.830
$h_{\mathrm{ss}}{=}1.01$ 108 13.5m 0.193 28.24 0.845
$h_{\mathrm{ss}}{=}1.10$* 107 13.5m 0.190 28.32 0.848
$h_{\mathrm{ss}}{=}1.20$ 99 13.6m 0.188 28.36 0.849
$h_{\mathrm{ss}}{=}1.30$ 100 13.7m 0.187 28.39 0.850
$h_{\mathrm{ss}}{=}1.50$ 92 13.8m 0.186 28.42 0.851
$h_{\mathrm{ss}}{=}2.00$ 75 14.2m 0.185 28.46 0.853
Table 7: Supersampling rate. Our rendering suffers from aliasing artifact so we render the image in $h_{\mathrm{ss}}\times$ higher resolution and apply image downsampling with anti-aliasing filter. The quality without supersampling is much worse than the others. Resampling the image with a very small $h_{\mathrm{ss}}=1.01$ can already boost quality significantly. We find the quality can keep going better with higher $h_{\mathrm{ss}}$ but the FPS drops by more than $30\%$ at $h_{\mathrm{ss}}=2$. More future development is needed for a more efficient anti-aliasing rendering of our method. We use $h_{\mathrm{ss}}=1.1$ for speed-quality trade-off.
[…]
$N_{\mathrm{shd}}{=}1$ 118 11.2m 0.201 27.43 0.840
$N_{\mathrm{shd}}{=}2$ 114 12.1m 0.193 27.94 0.847
$N_{\mathrm{shd}}{=}3$* 107 13.5m 0.190 28.32 0.848
Table 8: Degree of Spherical Harmonic (SH). The rendering time with higher SH degree is similar but the quality is much better. We use $N_{\mathrm{shd}}{=}3$ as our final setup. However, about $80\%$ of the parameters and the disk space is occupied by the SH coefficient with $N_{\mathrm{shd}}{=}3$. Future work may want to design a more parameters efficient representation for view-dependent colors.
[…]
$K{=}1$* 107 13.5m 0.190 28.32 0.848
$K{=}2$ 102 13.8m 0.189 28.32 0.849
$K{=}3$ 99 13.9m 0.189 28.33 0.849
Table 9: Number of sample points in a voxel when rendering. The effect of sampling more point inside a voxel is marginal as the voxels are typically subdivided into fine level with small size. It mainly affects the depth rendering for larger voxels. The figure shows the normal derived from the rendered depth. $K{=}1$ at the early training stage produce noisy depth as highlighted by the red arrow, while the depth noisy is mitigated when the voxels is subdivided into finer level. We suggest to use $K{>}1$ only when the sub voxel depth accuracy is required.
[…]
$h_{\mathrm{prune}}{=}0.01$ 96 16.7m 0.191 28.41 0.847
$h_{\mathrm{prune}}{=}0.03$ 107 13.5m 0.190 28.32 0.848
$h_{\mathrm{prune}}{=}0.05$* 119 11.6m 0.192 28.20 0.846
$h_{\mathrm{prune}}{=}0.10$ 158 9.1m 0.199 27.95 0.840
$h_{\mathrm{prune}}{=}0.15$ 188 7.7m 0.212 27.72 0.831
$h_{\mathrm{prune}}{=}0.20$ 213 6.8m 0.224 27.53 0.823
$h_{\mathrm{prune}}{=}0.30$ 241 5.9m 0.248 27.22 0.806
Table 10: Pruning threshold. We prune voxels with maximum blending weights below $h_{\mathrm{prune}}$. Higher FPS and faster processing time can be achieved by pruning more voxels but with loss in quality. We finally use $h_{\mathrm{prune}}{=}0.05$ to balance speed and quality.
[…]
$0.33$ 159 9.9m 0.210 27.98 0.833
$0.50$ 130 11.2m 0.197 28.16 0.843
$1.00$* 107 13.5m 0.190 28.32 0.848
$2.00$ 101 15.0m 0.190 28.36 0.848
$3.00$ 101 15.7m 0.190 28.36 0.848
Table 11: Subdivision scale. We subdivide $h_{\mathrm{percent}}{=}5$ percent of the voxels with the highest priority 15 times during the training. As the number of voxels become $(1+0.07h_{\mathrm{percent}})$ at each subdivision, the subdivision scales in above table shows their $\frac{(1+0.07h_{\mathrm{percent}})^{15}}{1.35^{15}}$, which indicate the expected relative number of voxels comparing to the base setup. The merit of subdividing more voxels each time is marginal comparing to the base setup.
[…]
$h_{\mathrm{ratio}}{=}1.0$ 120 11.6m 0.195 28.17 0.844
$h_{\mathrm{ratio}}{=}2.0$* 107 13.5m 0.190 28.32 0.848
$h_{\mathrm{ratio}}{=}3.0$ 103 14.7m 0.189 28.35 0.849
$h_{\mathrm{ratio}}{=}4.0$ 103 15.5m 0.189 28.38 0.849
Table 12: Initial ratio of the number of voxels in background and main regions. At the initialization stage, we heuristically subdivide voxel in the background region until the ratio of the number of voxel is $h_{\mathrm{ratio}}$ to the foreground region. The overall result quality are similar for different $h_{\mathrm{ratio}}$. It affects training time more than testing FPS as the training iterations per second before any pruning is depend on the initial number of voxels.
[…]
$\lambda_{\mathrm{tv}}{=}0$ 102 14.0m 0.202 27.77 0.832
$\lambda_{\mathrm{tv}}{=}1\text{e}{-}11$ 106 13.6m 0.196 27.97 0.840
$\lambda_{\mathrm{tv}}{=}1\text{e}{-}10$* 107 13.5m 0.190 28.32 0.848
$\lambda_{\mathrm{tv}}{=}1\text{e}{-}9$ 99 15.5m 0.213 27.85 0.822
Table 13: Total Variation (TV) loss. Similar to previous grid-based approaches [37, 11, 5], TV loss is also important in our method. We apply TV loss on density grid only for the first half $10{,}000$ iterations as applying TV for all iterations leads to blurrier rendering. TV with proper loss weighting leads to better quantitative results without loss of speed. The effect of TV loss is also visualized in above figure, where many geometric details emerge after the TV loss is turned off. The employed TV loss scheduling entails the coarse-to-fine optimization strategy.
[…]
$\lambda_{\mathrm{R}}{=}0$ 111 14.6m 0.200 28.11 0.843
$\lambda_{\mathrm{R}}{=}1\text{e}{-}4$ 110 14.6m 0.199 28.11 0.843
$\lambda_{\mathrm{R}}{=}1\text{e}{-}3$ 107 14.5m 0.196 28.20 0.845
$\lambda_{\mathrm{R}}{=}1\text{e}{-}2$* 107 13.5m 0.190 28.32 0.848
$\lambda_{\mathrm{R}}{=}1\text{e}{-}1$ 118 10.9m 0.205 27.77 0.830
Table 14: Color concentration loss. We find it helpful to apply L2 loss directly between observed pixel color and the individual voxel color of each voxel passing by the ray [37], which slightly improve training time and result quality.
[…]
$\lambda_{\mathrm{dist}}{=}0$ 105 14.9m 0.199 27.27 0.839
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}4$ 106 15.4m 0.199 27.48 0.839
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}3$ 105 15.1m 0.195 27.97 0.842
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}2$ 107 13.5m 0.190 28.32 0.848
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}1$ 137 9.9m 0.256 26.34 0.760
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}4$ from 10K 104 15.1m 0.199 27.40 0.839
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}3$ from 10K 105 15.0m 0.197 27.76 0.842
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}2$ from 10K 105 14.6m 0.193 28.08 0.845
$\lambda_{\mathrm{dist}}{=}1\text{e}{-}1$ from 10K* 113 13.7m 0.188 28.11 0.848
Table 15: Distortion loss. Distortion loss is proposed by MipNeRF-360 [2] and employed by many NeRF-based rendering approaches to encourage concentration of the blending weight distribution on a ray. We find distortion loss is also helpful in our method, especially for the PSNR. We also find that employing a larger distortion loss weight after the total variation loss is turned off lead to a cleaner geometry as shown in the above depth-derived normal visualization.
[…]
$\lambda_{\mathrm{T}}{=}1\text{e}{-}0$ 109 13.4m 0.192 28.23 0.847
$\lambda_{\mathrm{T}}{=}1\text{e}{-}3$ 109 13.5m 0.191 28.25 0.847
$\lambda_{\mathrm{T}}{=}1\text{e}{-}2$* 107 13.5m 0.190 28.32 0.848
$\lambda_{\mathrm{T}}{=}1\text{e}{-}1$ 109 12.8m 0.192 28.13 0.845
Table 16: Transmittance concentration loss. The effect of encouraging final ray transmittance to be either zero or one is marginal in the unbounded scenes. We find this loss is more important for the object-centric scenes with foreground region only and known background colors (e.g., Synthetic-NeRF dataset [27]).
[…]
neither* 107 13.5m 0.190 28.32 0.848
n-dmed 114 13.5m 0.191 28.10 0.849
n-dmean 97 14.0m 0.190 28.14 0.850
both 103 14.4m 0.191 27.99 0.849
Table 17: Mesh regularization losses for novel-view synthesis. We also try the normal-depth self-consistency losses for novel-view synthesis task. Despite of loss a little in PSNR, the regularization can make the rendered normals much smoother as shown in the visualization.
TnT dataset DTU dataset
$\mathcal{L}_{\mathrm{n\text{-}dmed}}$ $\mathcal{L}_{\mathrm{n\text{-}dmean}}$ $K$ F-$\uparrow$ Tr. $\downarrow$ Cf.$\downarrow$ Tr. $\downarrow$
3 0.56 10.1m 0.94 5.5m
3 0.59 10.1m 0.68 5.5m
3 0.61 10.6m 0.68 5.7m
1 0.61 10.7m 0.66 5.8m
2 0.61 10.8m 0.65 5.9m
3 0.62 10.9m 0.65 6.0m
Table 18: Mesh regularization losses. We show the results of the mesh regularization losses and the number of sample points when rendering a voxel on a subset of Tanks&Temples [20] and DTU [17] datasets.
E More Results
Results breakdown for novel-view synthesis.
In Tab. 19, we show details per-scene comparison with 3DGS [19] using our base setup. Our method uses much more primitives (i.e., voxels or Gaussians) compared to 3DGS on all the scenes. However, our average rendering FPS is still comparable to 3DGS. We find the FPS is scene-dependent, where we achieve much faster FPS on some of the scenes while slower on the others. Our method generally uses short training time. Regarding the quality metrics, our results are typically $-0.2$db PSNR and $-0.01$ SSIM behind 3DGS, while our LPIPS is better on average.
As discussed in the main paper, not only the scene representation itself affects the results, but the optimization and adaptive procedure are also an important factor. The strategy of adding more Gaussians progressively is not applicable to ours. We also have not explored to use of the coarse geometry estimated from SfM, while 3DGS uses SfM sparse points for initialization. As the first attempt of marrying rasterizer with fully explicit sparse voxels for scene reconstruction, there is still future potential for improvement from different aspects.
[…]
Scene 3DGS ours 3DGS ours 3DGS ours 3DGS ours 3DGS ours 3DGS ours
MipNeRF-360 [2] indoor scenes
bonsai 215 147 18.3 12.5 0.204 0.178 31.89 31.47 0.942 0.937 1.2M 6.1M
counter 160 102 20.7 14.6 0.199 0.184 29.03 28.62 0.909 0.896 1.2M 7.3M
kitchen 128 97 25.0 13.1 0.126 0.119 31.47 31.42 0.927 0.927 1.8M 7.9M
room 153 146 21.1 13.6 0.218 0.190 31.44 30.89 0.919 0.917 1.5M 8.8M
MipNeRF-360 [2] outdoor scenes
bicycle 72 164 31.9 10.8 0.211 0.206 25.18 24.93 0.765 0.755 6.1M 9.0M
garden 81 137 33.1 9.4 0.107 0.119 27.39 26.87 0.867 0.849 5.9M 8.7M
stump 110 142 25.5 11.2 0.216 0.219 26.61 26.11 0.772 0.756 4.9M 9.1M
treehill 123 170 22.4 11.8 0.327 0.284 22.47 22.24 0.632 0.609 3.7M 9.2M
flowers 137 125 22.0 12.1 0.335 0.277 21.57 21.40 0.606 0.617 3.6M 9.2M
DeepBlending [15] indoor scenes
drjohnson 116 261 25.0 7.8 0.244 0.258 29.11 29.11 0.901 0.887 3.3M 8.5M
playroom 163 260 19.7 7.4 0.244 0.228 30.08 30.04 0.907 0.893 2.3M 8.3M
Tanks&Temples [20] outdoor scenes
train 206 103 11.3 12.4 0.206 0.210 22.11 20.72 0.816 0.788 1.1M 9.3M
truck 154 117 16.3 10.3 0.147 0.119 25.40 24.74 0.882 0.873 2.6M 9.1M
Synthetic-NeRF [27] object scenes
chair 418 225 5.4 3.3 0.012 0.014 35.89 35.84 0.987 0.986 0.3M 2.5M
drums 406 275 5.8 2.9 0.037 0.044 26.16 26.10 0.955 0.946 0.3M 2.0M
ficus 476 399 5.0 2.1 0.012 0.015 34.85 34.35 0.987 0.984 0.3M 1.2M
hotdog 596 233 5.2 3.6 0.020 0.022 37.67 37.44 0.985 0.983 0.1M 2.6M
lego 415 175 5.9 4.4 0.015 0.017 35.77 35.38 0.983 0.980 0.3M 3.8M
materials 575 225 5.3 3.4 0.034 0.040 30.01 29.92 0.960 0.953 0.3M 2.5M
mic 344 361 5.6 2.0 0.006 0.008 35.38 35.72 0.991 0.991 0.3M 1.0M
ship 254 107 7.9 7.8 0.107 0.110 30.92 30.34 0.907 0.883 0.3M 5.9M
Table 19: Per-scene side-by-side comparison of our base setup with 3DGS [19].
Results breakdown for mesh reconstruction.
The F-score and chamfer distance of each scene from the Tanks&Temples and DTU [17] datasets are provided in Tab. 20. We only list the two representative NeRF-based methods and two GS-based methods in the result breakdown comparison. More methods with the average scores are in the main paper.
[…]
Method Barn Caterpillar Courthouse Ignatius Meetingroom Truck 24 37 40 55 63 65 69 83 97 105 106 110 114 118 122
NeuS 0.29 0.29 0.17 0.83 0.24 0.45 1.00 1.37 0.93 0.43 1.10 0.65 0.57 1.48 1.09 0.83 0.52 1.20 0.35 0.49 0.54
Neuralangelo 0.70 0.36 0.28 0.89 0.32 0.48 0.37 0.72 0.35 0.35 0.87 0.54 0.53 1.29 0.97 0.73 0.47 0.74 0.32 0.41 0.43
3DGS 0.13 0.08 0.09 0.04 0.01 0.19 2.14 1.53 2.08 1.68 3.49 2.21 1.43 2.07 2.22 1.75 1.79 2.55 1.53 1.52 1.50
2DGS 0.41 0.23 0.16 0.51 0.17 0.45 0.48 0.91 0.39 0.39 1.01 0.83 0.81 1.36 1.27 0.76 0.70 1.40 0.40 0.76 0.52
ours 0.35 0.33 0.29 0.69 0.19 0.54 0.61 0.74 0.41 0.36 0.93 0.75 0.94 1.33 1.40 0.61 0.63 1.19 0.43 0.57 0.44
Table 20: Result breakdown on Tanks&Temples [20] and DTU [17] datasets.
Synthetic dataset.
The results on the Synthetic-NeRF [27] dataset is provided in the last section of Tab. 19. We achieve good quality, high FPS, and fast training on this dataset. However, our quality is slightly worse than 3DGS [19] with slower FPS. Our development mainly focuses on real-world datasets. Future work may need more exploration to continue development on this dataset.
More qualitative results
We show qualitative comparison with 3DGS [19] on indoor and outdoor scenes in Fig. 11 and Fig. 12, respectively. Our visual quality is on par with 3DGS. We provide the visualization of the raw reconstructed meshes in Fig. 13. For quantitative evaluation, we follow previous works to apply mesh cleaning with the provided bounding box or masks. Despite the good quantitative results for meshes, some apparent artifacts can be observed from the visualization. In particular, our method focuses more on the geometric details and sometimes over-explains the texture on a flat surface with complex geometry. Future work may want to model a signed distance field instead of our current density field and introduce surface smoothness regularizers.
Figure 11: Qualitative novel-view rendering results on-par with 3DGS.
Figure 12: Qualitative novel-view rendering results on-par with 3DGS.
Figure 13: Qualitative results of the reconstructed mesh.
Acknowledgements.
We thank Min-Hung Chen and Yueh-Hua Wu for helpful paper flow suggestions and proofreading.